# A Generalized Information Bottleneck Theory of Deep Learning

### v2

@@ I.1 Sum-Versus-Whole Synergy
This alternative definition of GIB is based on the sum-versus-whole formulation of synergy, which compares information available from the complete feature set against the sum of information from individual components (Schneidman et al., 2003). The basic form is $\text{Syn}_{\text{GIB}}(\mathcal{X}\to Y)=I(\mathcal{X};Y)-\sum_{i=1}^{N}\left(I(X^{i};Y)\right)$, which considers only individual features. This captures the intuitive notion of synergy, for example, XOR has zero information from individual inputs but perfect information from their combination, yielding maximal synergy (Bell, 2003). Unlike exponentially complex PID-based measures (Williams & Beer, 2010), this formulation requires only $O(N)$ mutual information calculations, making it computationally feasible for tracking synergistic learning dynamics in high-dimensional neural networks . Combining this with our representation of the PMI-weighted combination of $Z$ and $Y$ and rewriting as a Lagrangian optimization we get the following:
$\mathcal{L}_{\text{SVW}}=\max_{p(Z|X)}\left[\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}I(\mathcal{X};Q(Z,Y))}}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{prediction term}}}-\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta^{-1}\sum_{i=1}^{N}I(X^{i};Q(Z,Y))}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{complexity term}}}\right]$ (22)

### v3

@@ I.1 Sum-Versus-Whole Synergy
This alternative definition of GIB is based on the sum-versus-whole formulation of synergy, which compares information available from the complete feature set against the sum of information from individual components (Schneidman et al., 2003). The basic form is $\text{Syn}_{\text{GIB}}(\mathcal{X}\to Y)=I(\mathcal{X};Y)-\sum_{i=1}^{N}\left(I(X^{i};Y)\right)$, which considers only individual features. This captures the intuitive notion of synergy, for example, XOR has zero information from individual inputs but perfect information from their combination, yielding maximal synergy (Bell, 2003). Unlike exponentially complex PID-based measures (Williams and Beer, 2010), this formulation requires only $O(N)$ mutual information calculations, making it computationally feasible for tracking synergistic learning dynamics in high-dimensional neural networks . Combining this with our representation of the PMI-weighted combination of $Z$ and $Y$ and rewriting as a Lagrangian optimization we get the following:
$\mathcal{L}_{\text{SVW}}=\max_{p(Z|X)}\left[\underbrace{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}I(Z;Y)}}_{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{prediction term}}}-\underbrace{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\beta^{-1}\sum_{i=1}^{N}I(X^{i};Q)}}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{complexity term}}}\right]$ (17)